# MicroScopiQ: Accelerating Foundational Models through Outlier-Aware Microscaling Quantization

### v3

@@ 7. Related Work
Model compression for LLMs. LLM weight Compression scheme like quantization often rely on mixed-precision quantization with different outlier and inlier bit-widths(Lee et al., 2024; Dettmers et al., 2023; Zhao et al., 2024; Jeong et al., 2024; Huang et al., 2024). For instance, SDQ (Jeong et al., 2024) decomposes a vector of weights into two separate inlier and outlier vectors each quantized at different precisions. Other class of quantization rely on channel-wise shared scaling with different scale factors for outliers (Lee et al., 2024; Shao et al., 2023; Huang et al., 2024). Other quantization techniques mitigate outliers from weights and activations (Liu et al., 2024c; Xiao et al., 2023; Ashkboos et al., 2024b) to make quantization easier.
LLM weight pruning methods (Frantar and Alistarh, 2023; Yin et al., 2023; Ashkboos et al., 2024a) rely on approximate solvers to identify the importance of weights based on magnitude of weights or weights activation combined (Sun et al., 2023). However, they fail to achieve satisfactory performance even at low sparsity levels(Lu et al., 2024). MicroScopiQ provides a unified solution for outlier-aware quantization by enabling flexible pruning of least-important weights to distribute additional outlier bits, thereby, simultaneously achieving high accuracy and hardware efficiency.

### v4

@@ 8. Related Work
Model compression for LLMs. LLM weight compression schemes like quantization often rely on mixed-precision quantization with different outlier and inlier bit-widths (Lee et al., 2024; Dettmers et al., 2023; Zhao et al., 2024; Jeong et al., 2024; Huang et al., 2024). Other class of quantization rely on channel-wise shared scaling with different scale factors for outliers (Lee et al., 2024; Shao et al., 2023; Huang et al., 2024). Other quantization techniques mitigate outliers from weights and activations (Liu et al., 2024d; Xiao et al., 2023; Ashkboos et al., 2024b).
LLM weight pruning methods (Frantar and Alistarh, 2023; Yin et al., 2023; Ashkboos et al., 2024a; Ramachandran et al., 2025a) rely on approximate solvers to identify the importance of weights based on magnitude of weights or weights activation combined (Sun et al., 2023). However, they fail to achieve satisfactory performance even at low sparsity levels(Lu et al., 2024). MicroScopiQ provides a unified solution for outlier-aware quantization by enabling flexible pruning of least-important weights to distribute additional outlier bits, thereby, simultaneously achieving high accuracy and hardware efficiency.
Unified pruning and quantization. SDQ (Jeong et al., 2024) combines pruning and quantization by decomposing weights into inlier and outlier vectors, each employing a fixed N:M pattern and different bit precisions. In contrast, MicroScopiQ removes the rigid N:M constraint and achieves hardware efficiency through a consistent bit-budget and unified data type per element. Moreover, unlike SDQ’s sequential pruning-then-quantization approach, MicroScopiQ integrates pruning implicitly within quantization, enabling tighter coupling.
@@
Figure 18. (a) Effect of time-multiplexed ReCoN units on the compute area and inference latency, (b) Implementation overhead of MicroScopiQ in NoC-based accelerators (Firoozshahian et al., 2023; Chen et al., 2019).
@@ 9. Conclusion
Acknowledgements.
This work was supported in part by CoCoSys, one of the seven centers in JUMP 2.0, a Semiconductor Research Corporation (SRC) program sponsored by DARPA. The authors would like to thank the anonymous reviewers for their valuable feedback and suggestions, which helped improve the quality of this paper. The authors also thank Jiawei Hu for place-and-route, Jianming Tong for help with designing ReCoN, and Zishen Wan for functional accuracy simulation and constructive suggestions.